THE UNIVERSITY OF CHICAGO 2008 Summer VIGRE Program for Undergraduates

Preference Form

0. Please give your name and program (Apprentice, SESAME, YSP).

1. Please indicate which of the course sequences seem of most interest to you, in order, using the numbering from the list below. (Note that the apprentice program courses are open to SESAME participants from the second week onwards, but conflict with YSP; all other courses are open to all participants).

2. Please indicate the general field of mathematics in which you feel most likely to write your paper for the program. You will be free to change your mind.

3. Any other information that might help us pair you with the right mentor?

LIST OF COURSES

0. Linear algebra and graph theory (Apprentice/SESAME; Abert and Babai)

1. Discrete mathematics (Babai)

2. Problem solving (Abert)

3. Geometry (Farb)

4. Probability (Lawler)

5. Topology (May)